Colony Inspection – 18th May

Observations:

weather improving and ghetto get warmer – 15C. Still below average. Wilson ref-queening may not have worked. Certainly old Aqua colony still intent on swarming. So decided mid-week to move Aqua to nuc and took to River Mole site. Aqua has 2 BIAS plus 2 frames of stores, and 1 of foundation. Put new colony (Blue) onto stand, still no sign of laying queen.

 

Flowering: Chesnut, wisteria and poppy

Hive Record
Colony
Q
BIAS
Brood
Stores
CommentNext Inspection
AquaNY3 BIAS frames, but also 3 charged QCs. About 3Kg in super3 in 80 in DT. Swarm prevention not worked.Will move queen to nuc
BlueNNNo BIAS, queen cell hatched outOKPlenty of beesCheck for laying queen
CyanYY6 + 7, not expanded2Kg in super, bringing in more honey than othersRe-arranged brood frames to add DTCheck for space
DenimNY7 + DT frameAbout 1 Kg in superLaying up DTRemark queen
EmeraldYY7 + 8 - expanded a bit, drone brood in top 1/21 Kg in super6 in 120 in DT, just OKKeep an eye out for swarming
FlaxYY6 BIAS, expanded so doing OKNot working super yetGoing OKRemove DT
GreenYY5 BIAS - not expandingOKStarting to lay up DTDecide if should be united - possibly with Aqua
HoneydewNY6 BIASAbout 1 Kg in superBees very ratty, flew up at me.Check for queen, not seen since strat of season